Lillich, Inc., manufactures and sells two products: Product U6 and Product R5. Data concerning the expected production of each product and the expected total direct labor-hours (DLHs) required to produce that output appear below:

Expected Production Direct Labor-Hours Per Unit Total Direct Labor-Hours
Product U6 750 9.5 7,125
Product R5 1,120 6.5 7,280
Total direct labor-hours 14,405

The direct labor rate is $27.60 per DLH. The direct materials cost per unit for each product is given below:

Direct Materials
Cost per Unit
Product U6 $250.40
Product R5 $167.80

The company is considering adopting an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product U6 Product R5 Total
Labor-related DLHs $ 201,638 7,125 7,280 14,405
Production orders orders 72,840 1,350 1,250 2,600
Order size MHs 1,020,608 6,500 6,800 13,300
$ 1,295,086

Which of the following statements concerning the unit product cost of Product U6 is true? (Round your intermediate calculations to 2 decimal places.)

rev: 03_25_2018_QC_CS-119201

  • The unit product cost of Product U6 under traditional costing is greater than its unit product cost under activity-based costing by $340.31.

  • The unit product cost of Product U6 under traditional costing is less than its unit product cost under activity-based costing by $5.63.

  • The unit product cost of Product U6 under traditional costing is greater than its unit product cost under activity-based costing by $5.63.

  • The unit product cost of Product U6 under traditional costing is less than its unit product cost under activity-based costing by $340.31.
Answer #1

1) Overhead rate per hour = 1295086/14405 = 89.91 per DLH

Unit product cost under traditional costing = 250.40+(9.5*27.60)+(9.5*89.91) = 1366.75

Calculate unit product cost under ABC Costing

Product U6
Labor related 99734
Production orders 37821
Order size 498793
Total overhead 636348
Unit 750
Overhead cost per unit 848.46
Direct material 250.40
Direct labor 262.20
Total 1361.06

The unit product cost of Product U6 under traditional costing is greater than its unit product cost under activity-based costing by $5.63.